# PHENIX Collaboration

### v1

@@ IV.2.2 The Coulomb interaction
$C_{2}(Q;\lambda,R,\alpha,N,\varepsilon)$ $=1-\lambda+\lambda C_{2}^{(0)}(Q;\lambda,R,\alpha,N,\varepsilon)\times\frac{\sum_{j}w_{j}K(q_{\rm inv})}{\sum_{j}w_{j}}$
${\rm with}\hskip 10.00002ptC_{2}^{(0)}(Q;R,\alpha,N,\varepsilon)$ $=(1+\exp(-R^{\alpha}Q^{\alpha}))\times N\times(1+\varepsilon Q),$ (9)
where $K(q_{\rm inv})$ is the Coulomb correction given by the parameterization and $w_{j}$ is the weight obtained from the two-dimensional actual pair distribution $A(Q,q_{\rm inv})$. The large $Q$ behavior was found to be consistent with a linear function, which is characterized with the functional form of $N\times(1+\varepsilon Q)$.
@@ VI Centrality and transverse-mass-dependent results
The transverse-mass ($m_{T}$) and centrality dependence (expressed in terms of the average number of participants, $N_{\rm part}$) of the Lévy parameters analyzed and are investigated along with their theoretically motivated combinations. The transverse mass is defined as $m_{T}~{}{=}~{}\sqrt{m_{\pi}^{2}+K_{T}^{2}}$, where $K_{T}=\sqrt{K_{x}^{2}+K_{y}^{2}}$ is the transverse component of the average momentum of the pair $K=0.5(p_{1}+p_{2})$ and $m_{\pi}$ is the pion mass. The number of participants $N_{\rm part}$ was determined via Glauber-model calculations by shown the PHENIX experiment based on Ref. [12]. The centrality dependencies of the Lévy parameters are characterized by theoretically or empirically motivated functions.
@@ VI.1 The centrality and transverse-mass dependence of the fit parameters
Figure 1: The transverse-mass dependence of the correlation-strength parameter $\lambda$ in six centrality bins. The central values are shown with dots, statistical uncertainties are indicated by vertical lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 2: The transverse-mass dependence of the correlation-strength parameter $\lambda$ in six centrality bins. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 3: The transverse-mass dependence of the Lévy-index parameter $\alpha$, shown in six centrality bins. It is parameterized with its $m_{T}$-averaged values $\alpha_{0}$. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 4: The transverse-mass dependence of the normalized correlation-strength parameter $\lambda/\lambda_{\rm max}$ for six centrality intervals. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties. For each centrality bin the data are fitted with the Gaussian function of Eq. (15). This parameterization can describe the data and is discussed in detail in Section VI.2. The fit parameters with the statistic and systematic uncertainties are shown in Fig. 7.
@@
Figure 5: The transverse-mass dependence of the inverse square of the Lévy-scale parameter $R$, shown in six centrality bins. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties. The values and uncertainties of the $A$ and $B$ linear fit parameters with the statistic and systematic uncertainties are shown in Fig. 8.
@@
The results of the fits for the 6 centrality and 23 $m_{T}$ bins are shown in Figs. 13. The intercept or the correlation-strength parameter $\lambda$ has a clearly observable suppression below $m_{T}\leq 0.5$ GeV average pair transverse mass. Above that it saturates at a centrality-dependent value. This low-$m_{T}$ suppression, which is characterized in more detail in Section VI.2 is observed in all centrality bins; however, in the most peripheral (50%–60%) centrality class, this observation starts to be limited by statistics. This is one of the reasons that no data are shown for the most peripheral, 60%–95% centrality class.
@@
The saturation value of the intercept parameter could depend on background processes which are not in the scope of the present analysis. As was shown in Ref. [33], the value of the intercept parameter is lower for Gaussian, intermediate for an Edgeworth expansion [23], and higher for a Lévy shape, which seems to be a systematic effect that depends on how the functional form of the Bose-Einstein correlation function extrapolates the data to the $Q=0$ limit. This systematic effect, an overall vertical uncertainty, can however be removed by normalizing to the saturation value of $\lambda(m_{T})$ at large $m_{T}$, following Refs. [33, 1]. The saturation value is taken as the average value of $\lambda(m_{T})$ in the interval 0.6 GeV $\leq{m_{T}}\leq$ 0.9 GeV and is denoted by $\lambda_{\rm max}$. This range is centrality independent. However, the value of $\lambda_{\rm max}$ does depend on centrality. The resulting $\lambda(m_{T})/\lambda_{\rm max}$ is shown in Fig. 4.
@@
Radial flow effects are known to be strongly centrality dependent, and hence are expected to significantly influence both $\lambda(m_{T})$ and $\lambda_{\rm max}$. Such an expectation is shown for example in Ref. [34]. If there is large radial flow, the decay products of $\eta^{\prime}$ are concentrated at high $m_{T}$ (compensating $p_{T}$ sharing between the several daughter pions, resulting in flat $\lambda(m_{T})/\lambda_{\rm max}(m_{T})$), while if the radial flow is low (such as in the peripheral event classes), then these decay products accumulate at low $m_{T}$ (resulting in a dip in $\lambda(m_{T})/\lambda_{\rm max}$ at low $m_{T}$). The two effects, the $p_{T}$ increasing radial flow and the $p_{T}$ sharing between the several pions from the $\eta^{\prime}\rightarrow\eta+\pi^{+}+\pi^{-}\rightarrow(\pi^{0}+\pi^{+}+\pi^{-})+\pi^{+}+\pi^{-}$ decay chain compensate one another at transverse velocity $\langle u_{T}\rangle{\approx}0.5$, as shown in Ref. [34]. This expectation is cross-checked with the Monte-Carlo simulations detailed in the Appendix, where the simulation-based centrality dependence of $\lambda(m_{T})/\lambda_{\rm max}$ (without any in-medium $\eta^{\prime}$ mass modification) is also shown.
@@
In Fig. 4, an approximate centrality independence of the characteristics of the suppression can be qualitatively observed. To test this observation quantitatively a $\chi^{2}$ test employs only statistical uncertainties to obtain conservative values of the CL. For each of the 15 possible pairs of centrality class, the $\chi^{2}$ is calculated under the hypothesis that their $\lambda/\lambda_{\rm max}$ distributions are identical. The resulting CL values are all in the range 0.15%–89.5%, which does not reject the hypothesis.
@@
The Lévy scale parameter, introduced in high energy particle and nuclear physics as the physical scale parameter $R$ in Ref. [17], decreases with $m_{T}$ for all centrality bins, as can be seen in Fig. 2. Analytic hydrodynamic calculations predict that $R^{-2}(m_{T})$ depends linearly on $m_{T}$, which can be parameterized as
@@
The measured values of the Lévy index $\alpha$ are shown in Fig. 3. In each centrality bin $\alpha$ appears to be independent of $m_{T}$. However, the $m_{T}$-averaged values, $\alpha_{0}$, which are indicated in Fig. 3, are found to depend significantly on centrality. The measured values of $\alpha$ are significantly lower than the Gaussian case of $\alpha=2$, and are significantly higher than the exponential/Cauchy case of $\alpha=1$. The centrality dependence of $\alpha_{0}$ is discussed in Section VI.2.
Similar to earlier PHENIX studies [35] in the 0%–30% centrality-class range, strong correlations are observed between the parameters $\lambda$, $R$, and $\alpha$. However, a specific combination of these three Lévy parameters, namely
@@
shown in Fig. 6, seems to be less correlated; hence it can be determined with smaller relative uncertainties. The correlation coefficients between the parameters $\lambda$, $R$ and $\alpha$ were determined in the Lévy fits and were considered when evaluating $\widehat{R}^{-1}$.
@@
Figure 6 shows a more detailed centrality dependent analysis. The centrality dependence of the $\widehat{A}$ slope parameter and $\widehat{B}$ intercept parameter is presented in Section VI.2. The linearity of $\widehat{R}^{-1}$ is broken at large $m_{T}$, which is likely due to the saturation of the $\lambda$ parameter in that range. As noted in Ref. [1], the understanding of the approximate scaling properties of $\widehat{R}$ is another currently unsolved theoretical challenge.
Figure 6: The transverse-mass dependence of the $\widehat{R}$ parameter, shown in six centrality bins. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties. The values and uncertainties of the linear fit parameters $\widehat{A}$ and $\widehat{B}$ with the statistic and systematic uncertainties are shown in Fig. 9.
@@ VI.2 The centrality dependence of the physical parameters
From the experimentally observed, and rather surprising, centrality-independent scaling of $\lambda(m_{T})/\lambda_{\rm max}$, was discussed in Section VI.1. Both the obtained $H$ and the $\sigma$ fit parameters that characterize the shape of $\lambda(m_{T})/\lambda_{\rm max}$ are consistent with the hypothesis of centrality independence.
Figure 7: The two parameters that characterize the suppression of the $\lambda/\lambda_{\rm max}(m_{T})$ and are defined in Eq. (15), (a) the magnitude ($H$), and (b) the width ($\sigma$) are shown as functions of $N_{\rm part}$. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties.
The affine linearity of the inverse square of the Lévy-scale parameter $R^{-2}(m_{T})$, Eq. (12), is demonstrated in Fig. 5. The parameters of these fits as a function of $N_{\rm part}$ are shown in Fig. 8. The slope parameter $A$ decreases with $N_{\rm part}$. The intercept parameter $B$, which could be connected to the size of the source, is slightly negative within statistical uncertainties but is compatible with zero if systematic uncertainties are taken into account. Note that in hydrodynamical calculations with $\alpha=2$, the parameter $B$ is dominated by the geometrical size of the whole source, while parameter $A$ is dominated by radial flow effects [36, 37, 38].
Figure 8: The parameters (a) $A$ and (b) $B$ of the affine linear fit to the inverse square of the Lévy-scale parameter that are defined in Eq.(12) as function of $N_{\rm part}$.
@@
Figure 9: The parameters (a) $\widehat{A}$ and (b) $\widehat{B}$ of the affine linear fit to $\widehat{R}$ that are defined in Eq. (14) as function of $N_{\rm part}$.
@@
Figure 10: The centrality dependence of $\alpha_{0}$, (a) the $m_{T}$-averaged value of $\alpha$ as a function of $N_{\rm part}$ and (b) the Lévy scale parameter $R$ as a function of $N_{\rm part}^{1/3}$ in selected $m_{T}$ bins.

### v2

@@ IV.2.2 The Coulomb interaction
$C_{2}(Q;\lambda,R,\alpha,N,\varepsilon)$ $=1-\lambda+\lambda\>C_{2}^{(0)}(Q;R,\alpha,N,\varepsilon)\>w(Q;R,\alpha)$
${\rm with}\hskip 10.00002ptC_{2}^{(0)}(Q;R,\alpha,N,\varepsilon)$ $=(1+\exp(-R^{\alpha}Q^{\alpha}))N(1+\varepsilon Q)$
${\rm and}\hskip 10.00002ptw(Q;R,\alpha)$ $=\frac{\sum_{k}A(Q,q_{{\rm inv,}k})K(q_{{\rm inv,}k};R,\alpha)}{\sum_{k}A(Q,q_{{\rm inv,}k})}$ (9)
where $K(q_{{\rm inv,}k};R,\alpha)$ is the Coulomb correction given by the parameterization [32, 33]. In the definition of the weight function $w(Q;R,\alpha)$, the index of summation $k$ runs over those bins in $q_{\rm inv}$ for a given value of $Q$, where the number of actual pairs after the two-track cuts is nonvanishing. This summation thus averages over the $q_{\rm inv}$-dependent Coulomb correction for a Lévy-type source characterized by $R$ and $\alpha$, in $q_{\rm inv}$ bins only where $A(Q,q_{{\rm inv,}k})$ is nonzero. This method generalizes on Eqs. (2) and (3) of Ref. [21] for a Lévy-shaped Bose-Einstein correlation function and also accounts for the $q_{\textmd{inv}}$ dependence of the Coulomb correction. The large $Q$ behavior was found to be consistent with a linear function, which is characterized with the functional form of $N(1+\varepsilon Q)$.
@@ VI Centrality and transverse-mass-dependent results
[…]
@@ VI.1 The centrality and transverse-mass dependence of the fit parameters
Figure 1: The transverse-mass dependence of the correlation-strength parameter $\lambda$ in six centrality bins obtained from Lévy fits with Eq. (9). The central values are shown with dots, statistical uncertainties are indicated by vertical lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 2: The transverse-mass dependence of the Lévy-scale parameter $R$ in six centrality bins obtained from Lévy fits with Eq. (9). The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 3: The transverse-mass dependence of the Lévy-index of stability parameter $\alpha$, shown in six centrality bins obtained from Lévy fits with Eq. (9). The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 4: The transverse-mass dependence of the normalized correlation-strength parameter $\lambda/\lambda_{\rm max}$ for six centrality intervals obtained by rescaling Fig. 1 with a centrality-dependent $\lambda_{\textmd{max}}$ defined as the average value of $\lambda(m_{T})$ in the $0.45\leq m_{T}\leq 0.9$ GeV interval. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties. For each centrality bin the data are fitted with the Gaussian function of Eq. (15). This parameterization can describe the data and is discussed in detail in Section VI.2. The fit parameters with the statistic and systematic uncertainties are shown in Fig. 7.
@@
Figure 5: The transverse-mass dependence of the inverse square of the Lévy-scale parameter $R$, shown in six centrality bins. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties. For each centrality bin the data are fitted with the linear function of Eq. (12). This parameterization can describe the data and is discussed in detail in Section VI.2. The fit parameters with the statistic and systematic uncertainties are shown in Fig. 8.
@@
The results of the fits for the 6 centrality and 23 $m_{T}$ bins are shown in Figs. 13. The intercept or the correlation-strength parameter $\lambda$ has a clearly observable suppression below $m_{T}\leq 0.5$ GeV average pair transverse mass. Above that it saturates at a centrality-dependent value. This low-$m_{T}$ suppression, which is characterized in more detail in Section VI.2 is observed in all centrality bins; however, in the most-peripheral (50%–60%) centrality class, this observation starts to be limited by statistics. This is one of the reasons that no data are shown for the most peripheral, 60%–95% centrality class.
@@
The saturation value of the intercept parameter could depend on background processes which are not in the scope of the present analysis. As was shown in Ref. [35], the value of the intercept parameter is lower for Gaussian, intermediate for an Edgeworth expansion [25], and higher for an exponential shape, which seems to be a systematic effect that depends on how the functional form of the Bose-Einstein correlation function extrapolates the data to the $Q=0$ limit. This systematic effect, an overall vertical uncertainty, can however be removed by normalizing to the saturation value of $\lambda(m_{T})$ at large $m_{T}$, following Refs. [35, 1]. The saturation value is taken as the average value of $\lambda(m_{T})$ in the interval 0.45 GeV $\leq{m_{T}}\leq$ 0.9 GeV and is denoted by $\lambda_{\rm max}$. This centrality-independent range is the same for each of the six centrality classes considered in this manuscript. However, from the similar range considered in Ref. [1] for the 0%–30% centrality selection, this range is slightly modified due to the different centrality classes considered. This range modification shifts the central values of $\lambda_{\textmd{max}}$ slightly, but the modification is within one standard deviation, which is within the uncertainties given in Ref. [1]. However, the value of $\lambda_{\rm max}$ does depend on centrality. The resulting $\lambda(m_{T})/\lambda_{\rm max}$ ratio is shown in Fig. 4.
@@
Radial-flow effects are known to be strongly centrality dependent, and hence are expected to significantly influence both $\lambda(m_{T})$ and $\lambda_{\rm max}$. Such an expectation is shown for example in Ref. [36]. If there is large radial flow, the decay products of $\eta^{\prime}$ are concentrated at high $m_{T}$ (compensating $p_{T}$ sharing between the several daughter pions, resulting in flat $\lambda(m_{T})/\lambda_{\rm max}(m_{T})$), while if the radial flow is low (such as in the peripheral event classes), then these decay products accumulate at low $m_{T}$ (resulting in a dip in $\lambda(m_{T})/\lambda_{\rm max}$ at low $m_{T}$). The two effects, the $p_{T}$ increasing radial flow and the $p_{T}$ sharing between the several pions from the $\eta^{\prime}\rightarrow\eta+\pi^{+}+\pi^{-}\rightarrow(\pi^{0}+\pi^{+}+\pi^{-})+\pi^{+}+\pi^{-}$ decay chain compensate one another at transverse velocity $\langle u_{T}\rangle{\approx}0.5$, as shown in Ref. [36]. This expectation is cross-checked with the Monte-Carlo simulations detailed in the Appendix, where the simulation-based centrality dependence of $\lambda(m_{T})/\lambda_{\rm max}$ (without any in-medium $\eta^{\prime}$ mass modification) is also shown.
@@
In Fig. 4, an approximate centrality independence of the characteristics of the suppression can be qualitatively observed. To test this observation quantitatively, a $\chi^{2}$ test was employed to obtain conservative values of the CL using only statistical uncertainties. For each of the 15 possible pairs of centrality classes, the $\chi^{2}$ is calculated under the hypothesis that their $\lambda/\lambda_{\rm max}$ distributions are identical. The resulting CL values are all in the range 0.15%–89.5%, which does not reject the hypothesis.
@@
The Lévy-scale parameter, introduced in high-energy particle and nuclear physics as the physical-scale parameter $R$ in Ref. [10], decreases with $m_{T}$ for all centrality bins, as can be seen in Fig. 2. Analytic hydrodynamic calculations [27, 37, 38] predict that in the $\alpha=2$ special case, $R^{-2}(m_{T})$ depends linearly on $m_{T}$, which can be parameterized as
@@
Figure 3 shows the measured values of the Lévy index $\alpha$. In each centrality bin $\alpha$ appears to be independent of $m_{T}$. In each centrality class, the data of Fig. 3 can be well represented by the $m_{T}$-averaged values, $\alpha_{0}$. As indicated in Fig. 10, these values are found to depend on centrality. The measured values of $\alpha$ are significantly lower than the Gaussian case of $\alpha=2$, and are significantly higher than the exponential/Cauchy case of $\alpha=1$. The centrality dependence of $\alpha_{0}$ is discussed in Section VI.2.
Similar to earlier PHENIX studies in the 0%–30% centrality-class range  [1], strong correlations are observed between the parameters $\lambda$, $R$, and $\alpha$. However, a specific combination of these three Lévy parameters, namely
@@
shown in Fig. 6, seems to be less correlated with the direct fit of Lévy parameters $\lambda$, $R$, $\alpha$ as compared to the correlations among them, which is very similar to the correlation plots published in Ref. [1] for the 0%–30% centrality class. The correlation coefficients between the parameters $\lambda$, $R$ and $\alpha$ were determined in the Lévy fits and were considered when evaluating $\widehat{R}^{-1}$.
@@
Figure 6 shows a more detailed centrality-dependent analysis. The centrality dependence of the $\widehat{A}$ slope parameter and $\widehat{B}$ intercept parameter is presented in Section VI.2. The linearity of $\widehat{R}^{-1}$ is broken at large $m_{T}$, which is likely due to the saturation of the $\lambda$ parameter in that range. As noted in Ref. [1], the understanding of the approximate scaling properties of $\widehat{R}$ is another currently unsolved theoretical challenge.
Figure 6: The transverse-mass dependence of the $\widehat{R}$ parameter, shown in six centrality bins. The central values are shown with dots, statistical uncertainties are indicated by vertical black lines, while boxes are used to illustrate the systematic uncertainties. For each centrality bin the data are fitted with the linear function of Eq. (14). This parameterization can describe the data and is discussed in detail in Section VI.2. The fit parameters with the statistic and systematic uncertainties are shown in Fig. 9.
@@ VI.2 The centrality dependence of the physical parameters
From the experimentally observed, and rather surprising, centrality-independent scaling of $\lambda(m_{T})/\lambda_{\rm max}$, the centrality independence of the $H$ and the $\sigma$ fit parameters is expected, as discussed in Section VI.1. Indeed, both obtained parameters are consistent with the hypothesis of centrality independent values $H_{0}$ and $\sigma_{0}$, as indicated in Fig. 7.
Figure 7: The two parameters that characterize the suppression of the $\lambda/\lambda_{\rm max}(m_{T})$ and are defined in Eq. (15), (a) the magnitude ($H$), and (b) the width ($\sigma$) are shown as functions of $N_{\rm part}$. The central values are shown with dots, statistical uncertainties are indicated by vertical [red] lines, while boxes are used to illustrate the systematic uncertainties. Both panels (a) and (b) are consistent with a centrality-independent constant values $H_{0}$, $\sigma_{0}$.
The affine linearity of the inverse square of the Lévy-scale parameter $R^{-2}(m_{T})$, Eq. (12), is demonstrated in Fig. 5. The parameters of these fits as a function of $N_{\rm part}$ are shown in Fig. 8. The slope parameter $A$ decreases with $N_{\rm part}$. The intercept parameter $B$, which could be connected to the size of the source, is slightly negative within statistical uncertainties but is compatible with zero if systematic uncertainties are taken into account. Note that in hydrodynamical calculations with $\alpha=2$, the parameter $B$ is dominated by the inverse of the squared geometrical size of the whole source: the smaller the value of B, the larger the geometrical size of the source, while parameter $A$ is dominated by radial flow effects [39, 38, 40].
Figure 8: The two parameters of the affine linear fit to the inverse square of the Lévy-scale parameter $1/R^{2}(m_{T})$ are defined in Eq.(12), (a) the slope parameter ($A$), and (b) the intercept parameter ($B$) are shown as functions of $N_{\rm part}$. The central values are shown with dots, statistical uncertainties are indicated by vertical [red] lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 9: The two parameters of the affine linear fit to the inverse square of the Lévy-scale parameter $1/\widehat{R}(m_{T})$ are defined in Eq.(14), (a) the slope parameter ($\widehat{A}$), and (b) the intercept parameter ($\widehat{B}$) are shown as functions of $N_{\rm part}$. The central values are shown with dots, statistical uncertainties are indicated by vertical red lines, while boxes are used to illustrate the systematic uncertainties.
@@
Figure 10: The centrality dependence of (a) $\alpha_{0}$, the $m_{T}$-averaged value of $\alpha$ as a function of $N_{\rm part}$ and (b) the Lévy scale parameter $R$ as a function of $N_{\rm part}^{1/3}$ in selected $m_{T}$ bins.